Find the covariance of the returns of the two assets below. These are historical returns so each outcome is equally likely.

          Year             Asset A                  Asset B

          2007             30.0%                     15.0%

          2008             - 8.0%                     22.7%

          2009             2.0%                     -14%

          2010             7.5%                     2.8%

          2010             1.0%                      12.0%
Answer #1

Let Returns for A be denoted by X and For B by Y

Average return of A = x = (0.30 - 0.08 + 0.02 + 0.075 + 0.01)/5 = 0.065

Average return of B = y = (0.15 + 0.227 - 0.14 + 0.028 + 0.12)/5 = 0.077

Covariance = Σ (X - x)(Y - y) / N

= [ (0.30-0.065)(0.15-0.077) + (-0.08-0.065)(0.227-0.077) + (0.02-0.065)(-0.14-0.077) + (0.075-0.065)(0.028-0.077) + (0.01-0.065)(0.12-0.077) ]/5

= 0.000463